OpenStudy (anonymous):

If x and y vary directly, and x = 10 when y = 3, what will be the value of x when y = 9?

OpenStudy (hayleymeyer):

from x = 10 and y = 3 we have: 3 = k*10 so k = 3/10 and we have: y = (3/10)x evaluating at y = 9: 9 = (3/10)x x = (10/3)*9 x = 30

OpenStudy (anonymous):

go hayhay

OpenStudy (hayleymeyer):

Bam!!

OpenStudy (anonymous):

thank you!!